OpenStudy (antonio_xx2):

Match each shape with its surface area. Use three point one four for pi and round to the nearest whole number where necessary. 1.square pyramid: base length 6 meters, slant height 10 meters 2.cone: radius 4 meters, slant height 8 meters 3.square pyramid: base length 5 meters, slant height 11 meters 4.cone: radius 6 meters, slant height 4 meters A)1511m^2 B)156m^2 C)188m^2 D)135m^2

OpenStudy (texaschic101):

surface area of square pyramid : 2bs + b^2 b = 6 s = 10 now we sub 2(6)(10) + 6^2 2(60) + 36 120 + 36 156 m^2 surface area of cone : (pi )rs + (pi) r^2 r = 4 s = 8 pi = 3.14 now we sub (3.14)(4)(8) + (3.14)(4^2) (3.14)(32) + (3.14)(16) 100.48 + 50.24 150.72 square pyramid : 2bs + b^2 b = 5 s = 11 now we sub 2(5)(11) + 5^2 2(55) + 25 110 + 25 135 m^2 cone : (pi) rs + (pi) r^2 r = 6 s = 4 pi = 3.14 now we sub (3.14)(6)(4) + (3.14)(6^2) 3.14(24) + 3.14(36) 75.36 + 113.04 188.4 m^2 I am not getting the exact same answers for the cones

OpenStudy (antonio_xx2):

thanks:)